Determine for which value of "\alpha" the given series converge absolutely, converge conditionally, or diverges.

  1. "\displaystyle\sum_{k=1}^ ∞ ((-1)" k+1 * "\alpha^k" ) / k2
  2. "\\displaystyle\\sum_{k=1}^ \u221e ((\\alpha -1)^k\n\n\u200b" ) / k

For each of the following series, determine whether it is absolutely convergent, conditionally convergent, or divergent. Justify your answer.

  1. "\displaystyle\sum_{n=2}^ ∞ (-1)^n/(n ln n)"
1
Expert's answer
2021-11-23T16:46:06-0500

1.

series diverges if

"\displaystyle{\lim_{k\to \infin}}|a_k|=\displaystyle{\lim_{k\to \infin}}|\alpha^k/k^2|=\infin\implies |\alpha|>1"


series converge absolutely if

"\displaystyle{\lim_{k\to \infin}}|a_k|=\displaystyle{\lim_{k\to \infin}}|\alpha^k/k^2|=0\implies |\alpha|<1"


2.

series diverges if

"\displaystyle{\lim_{k\to \infin}}|a_k|=\displaystyle{\lim_{k\to \infin}}|(\alpha-1)^k/k|=\infin\implies |\alpha|>1"


series converge absolutely if

"\displaystyle{\lim_{k\to \infin}}|a_k|=\displaystyle{\lim_{k\to \infin}}|(\alpha-1)^k/k|=0\implies |\alpha-1|<1\implies 0<\alpha<1" and "1<\alpha<2"


3.

for alternating series:

series converges if

"|a_n|" decreases monotonically, i. e. "|a_{n+1}|\le |a_n|" and "\displaystyle{\lim_{n\to \infin}}|a_n|=0"


"\displaystyle{\lim_{n\to \infin}}|a_n|=\displaystyle{\lim_{n\to \infin}}|1/(nlnn)|=0"


"\frac{1}{nlnn}\ge\frac{1}{(n+1)ln(n+1)}"


so, series converge


for series of absolute values "\displaystyle\sum_{n=2}^{\infin}\frac{1}{nlnn}" :


"\int_2^{\infin}\frac{1}{xlnx}dx=ln(lnx)|^{\infin}_2=\infin"


so, "\displaystyle\sum_{n=2}^{\infin}\frac{1}{nlnn}" diverges, this means that "\displaystyle\sum_{n=2}^{\infin}\frac{(-1)^n}{nlnn}" converge conditionally